# LTOS: Layout-controllable Text-object Synthesis via Adaptive Cross-attention Fusions

### v2

@@ III Experiment
Experimental settings
Our backbone is based on GLIGEN [23]. We initialize visual-text rendering module with pre-trained “bounding box + text” weights. Our dataset consists of 217,820 training and 11,113 testing samples. For diffusion, we set the downsampling factor $f=8$, and the latent dimension is 64$\times$64$\times$4. We train the model for 1.3 M iterations with batch size 24. Since LTOS is a novel designed task, we mainly compare our model to the following SOTA baselines: including layout-to-image generation model GLIGEN [23]${}_{\text{CVPR'22}}$. We compare our model to text rendering baselines with TextDiffuser [31]${}_{\text{NeurIPS'23}}$, AnyText [6]${}_{\text{ICLR'24}}$ and GlyphControl [7]${}_{\text{NeurIPS'23}}$. We use three metrics: 1) OCR. Acc [22], the word-level accuracy of rendered texts; 2) Normalized Edit Distance (NED) [32], the similarity between two strings; 3) AP [33], the accuracy of object generation.
Main results
As Table I, our model significantly outperform all state-of-the-art baselines on the benchmark dataset in all metrics (+7.52% on OCR. ACC and +9.21% on NED). AnyText performs bad in OCR. ACC and NED, since it hardly automatically wraps and many multi-line texts in the LTOS dataset. Because the text rendered by GlyphControl is only rotated up to 20 degrees clockwise or counterclockwise, we fine-tuned on the LOST dataset, the results were still not ideal compared to our model.
Ablation studies
In Table II, removing any of the modules results in worse results. Row 1 and 4 verify the effectiveness of text-object self-adaptive fusion. Row 2 ascertains the validity of the self-adaptive mechanism in Sec. II-D. Row 3 in Table II) suggests that text perceptual loss enhances the quality of text generation.
Analyses on attention-layers
We test the results with different numbers and settings of injecting cross-attention layers (4-layer Dense: four layers after the 0-th to 3-rd zero-convolution layers; 4-layer Sparse: four layers after the 0-th, 3-rd, 6-th and 9-th zero-convolution layers) in Table I. It suggests that increasing the number of layers doesn’t necessarily lead to better results. Instead, over-introduction of layers can lead to significant results degradation. Results show the 4-layer Dense leads to best results, which is used in our paper.
TABLE I: TOF’s resuls compared to baselines and analysis of fusion cross-attention layers. * denotes finetuning on LTOS dataset. We report OCR. ACC, NED and AP scores.
@@
TextDiffuser [15] 0.2551 0.4824 -
$\text{GlyphControl}^{*}$ [7] 0.3601 0.5404 -
AnyText [6] 0.1732 0.3274 -
GLIGEN [23] 0.0018 0.0092 0.4958
@@
Figure 3: Qualitative comparison of TOF and the baselines. Due to the certain differences between our input and baselines, we keep the input to be the same based on the input format accepted by all models and implementation details.
Qualitative results
Our qualitative results compared with baselines and widely used text-to-image generation APIs (SD-XL, Midjourney, DeepFloyd IF and Bing Image Creator) are visualized in Fig. 3 (a). It shows that our method renders more legible, clearer visual texts and plausible objects at the predefined positions. Our model supports arbitrary rotation of text area, tiny font and different thicknesses of font in Fig.  3 (b). We observed notable limitations in the baselines. AnyText cannot automatically wrap resulting in garbled text. TextDiffuser and GlyphControl fail to render tiny-font text and some special characters. Further, they all sometimes generate objects of the wrong categories or omit objects.
@@ IV CONCLUSION
We define an unified task LTOS accurately controlling visual text and object generation and propose the LTOS dataset. Further, we propose a novel TOF model with a text-object self-adaptive cross-attention fusion module balancing text and object synthesis. Experiments show that TOF excels SOTA baselines for text rendering maintaining comparable results to the object generation model.

### v3

@@ III Experiment
Settings. Our backbone is based on GLIGEN [li2023gligen]. We initialize the visual-text rendering with pre-trained “bounding box + text” weights. Our dataset has 217,820 training and 11,113 testing samples. For diffusion, we set the downsampling factor $f=8$ and the latent dimension is 64$\times$64$\times$4. We train the model for 1.3 M iterations with batch size 24. Since LTOS is a new task, we compare with baselines on related tasks: (1) layout-to-image generation tasks, GLIGEN [li2023gligen], (2) text rendering tasks, TextDiffuser [zhang2024brush], AnyText [tuo2023anytext], and GlyphControl [yang2024glyphcontrol] (SOTA on MARIO-10M, AnyWord-3M, LAION-Glyph datasets, respectively). We use three metrics: 1) OCR. Acc [li2022pp], the word-level accuracy of rendered texts; 2) Normalized Edit Distance (NED) [Marzal1993pami],the similarity between two strings; 3) AP [li2021image], the accuracy of object generation.
Main results. As Tb I, our model significantly outperform all state-of-the-art baselines on the benchmarks in all metrics (+7.10% on OCR. ACC and +6.63% on NED). TextDiffuser generates text layout which is not effective at rendering non-center text. When AnyText renders multiple texts with different fonts, sizes and rotation angles on backgrounds with different textures, text’s clarity is affected.
Notice that GlyphControl performs poorly when rendering rotated text beyond 20 degrees especially on complex background textures.
Ablation studies. In Tb II, removing any of the modules results in worse results. Row 2 and 5 verify the effectiveness of text-object self-adaptive fusion. Row 3 verify the self-adaptive mechanism in Sec. II-D. Row 4 in Tb II suggests that text perceptual loss enhances the quality of text generation.
Analyses on attention-layers. We test the model with different numbers and settings of injecting cross-attention layers (4-layer Dense: four layers after the 0-th to 3-rd zero-convolution layers; 4-layer Sparse: four layers after the 0-th, 3-rd, 6-th and 9-th zero-convolution layers) in Tb I. It suggests that increasing the number of layers doesn’t necessarily lead to better results. Instead, the over-introduction of layers leads to significant results degradation. We use the 4-layer Dense setting in this paper, which leads to the best results.
TABLE I: TOF compared to baselines and analysis on cross-attention. * means finetuning on LTOS dataset. - indicates metrics do not fit baselines considering the baselines’ original tasks.
@@
TextDiffuser [chen2024textdiffuser] 0.2551 0.4824 -
$\text{GlyphControl}^{*}$ [yang2024glyphcontrol] 0.3643 0.5662 -
AnyText [tuo2023anytext] 0.2862 0.5074 -
GLIGEN [li2023gligen] - - 0.4958
@@
Figure 3: Qualitative comparison of TOF and the baselines. Due to the differences between our input and baselines, we keep the input to be the same based on the input format accepted by all models.
Qualitative results. Our qualitative results compared with baselines and widely used T2I generation APIs (SD-XL, Midjourney, DeepFloyd IF, and Bing Image Creator) are visualized in Fig. 3 (a). It shows that our method renders more legible, clearer visual texts and plausible objects at the predefined positions. Our model supports arbitrary rotation of text area, tiny font, and different thicknesses of the font in Fig.  3 (b). We observed limitations in baselines. AnyText can’t easily wrap and rotate text at special angles resulting in garbled text. TextDiffuser and GlyphControl fail to render tiny-font text and some special characters. They all sometimes generate objects of the wrong categories or omit objects.
@@ IV CONCLUSION
We define a unified task LTOS accurately controlling visual text and object generation and propose the LTOS dataset. We propose a TOF model with a text-object self-adaptive cross-attention fusion module balancing text and object synthesis. Experiments show that TOF obtains SOTA on LTOS and performs well in text rendering and object generations.